# AraHealthQA 2025: The First Shared Task on Arabic Health Question Answering

### v2

@@ Abstract
We introduce AraHealthQA 2025, the Comprehensive Arabic Health Question Answering Shared Task, held in conjunction with ArabicNLP 2025 (co-located with EMNLP 2025). This shared task addresses the paucity of high-quality Arabic medical QA resources by offering two complementary tracks: MentalQA, focusing on Arabic mental health Q&A (e.g., anxiety, depression, stigma reduction), and MedArabiQ, covering broader medical domains such as internal medicine, pediatrics, and clinical decision making. Each track comprises multiple subtasks, evaluation datasets, and standardized metrics, facilitating fair benchmarking. The task was structured to promote modeling under realistic, multilingual, and culturally nuanced healthcare contexts. We outline the dataset creation, task design and evaluation framework, participation statistics, baseline systems, and summarize the overall outcomes. We conclude with reflections on the performance trends observed and prospects for future iterations in Arabic health QA.
AraHealthQA 2025: The First Shared Task on Arabic Health Question Answering
Hassan Alhuzali1, Farah Shamout2, Muhammad Abdul-Mageed3, Chaimae Abouzahir2, Mouath Abu-Daoud2, Ashwag Alasmari4, Walid Al-Eisawi2, Renad Al-Monef 4, Ali Alqahtani4, Lama Ayash4, Nizar Habash2, Leen Kharouf2, 1Umm Al-Qura University, 2New York University Abu Dhabi, 3The University of British Columbia, 4King Khalid University, Correspondence: hrhuzali@uqu.edu.sa & farah.shamout@nyu.edu
@@ 2.1 Mental Health Benchmarks
To bridge this gap, the MentalQA dataset Alhuzali and Alasmari (2025); Alhuzali et al. (2024) was developed, providing annotated Arabic question-answer pairs that cover a variety of question types and answer strategies. This dataset supports the creation and evaluation of NLP systems capable of handling diverse mental health inquiries, forming the foundation of the Arabic Mental Health Question Answering Shared Task. By leveraging the MentalQA dataset, this track provides a dedicated benchmark for Arabic mental health question-answering. It addresses the dual challenge of classification and response generation, creating a platform to systematically evaluate models in a culturally sensitive setting. Through this effort, the tack of MentalQA promotes research on the building of reliable and context-sensitive NLP systems for Arabic-speaking communities.
@@ 3.1.1 Subtask 1 and 2
For both subtasks, the dataset is based on MentalQA and is divided into 300 samples for training, 50 samples for development, and 150 samples for testing. The training set can be used to fine-tune LLMs or serve as a base for few-shot learning approaches. The development set is intended to tune hyper-parameters and evaluate performance, while the test set ensures fair benchmarking of all participants.
@@ 4 Shared Task Teams
Submission Rules: For Track 1, we allowed participant teams to submit up to five runs for each test set and for each of the three subtasks. For Track 2, participants were initially allowed 10 submissions each, which was later increased to 15 submissions due to platform-specific issues. For each team, only the submission with the highest score was retained on the official leaderboard. The official evaluation relied on a blind test set. To ensure fairness and reproducibility, each subtask of each track was hosted as a separate competition on Codabench, enabling automatic scoring and ranking of submissions. These Codabench instances will remain active even after the official competition concludes, supporting continued experimentation and benchmarking on the MentalQA and MedArabiQ datasets.
@@ 4.1 Track 1: MentalQA
Evaluation: Subtask 1 and Subtask 2 are multi-label classification tasks and are evaluated using Weighted F1 score and Jaccard score. The Weighted F1 balances precision and recall while accounting for class imbalance, whereas the Jaccard score measures the overlap between predicted and gold label sets, making it suitable for multi-label evaluation. Subtask 3 is evaluated using BERTScore, which leverages contextual embeddings from pre-trained language models to capture semantic similarity between generated responses and reference answers. Together, these metrics provide a robust assessment of system performance across the classification and generation subtasks, reflecting both the accuracy of the labels and the semantic quality of the outputs.
@@
MarsadLab Biswas et al. (2025) Hamad Bin Khalifa University, Qatar; Northwestern University, Qatar 1, 2, 3
Sindbad Zirikly et al. George Washington University, USA 1, 2, 3
@@
RetAug Amr et al. (2025) Nile University, Egypt 1, 3
@@
Fahmni Rasmy et al. (2025) MBZUAI, UAE; Gameball Company; German International University, Egypt; American University in Cairo, Egypt 1, 2, 3
@@ 5.1.4 General Description of Submitted Systems (Track 1)
Subtask 1: The winning team, mucAI, achieved a weighted F1-score of 0.61 for question classification. Their system, Explain–Retrieve–Verify (ERV), is a lightweight, training-free pipeline for multi-label categorization of Arabic mental-health questions. ERV combines a chain-of-thought LLM classifier with example-based retrieval and a verification agent. The LLM proposes candidate labels and rationales, a similarity agent retrieves top-$k$ nearest questions via multilingual sentence-transformer embeddings to provide case-based priors, and the verification agent reconciles these signals to produce a final label set with calibrated confidence. A post-processing step handles code parsing and confidence clamping. ERV runs efficiently at inference time without requiring fine-tuning or external data.
Subtask 2: The winning team, Sindbad, achieved a weighted F1-score of 0.71 and a Jaccard score of 0.71 for answer classification. Their approach leverages dataset augmentation to balance underrepresented classes, followed by a rigorous pipeline that uses state-of-the-art pre-trained language models (PLMs) and large language models (LLMs) for few-shot prompting and instruction fine-tuning. They utilize Gradient-free Edit-based Instruction Search (GrIPS) to optimize prompt selection, improving the quality and consistency of the QA system without extensive manual intervention.
Subtask 3: The winning team, RetAug, achieved a BERTScore of 0.679 for generative question answering. Their system employs a Retrieval-Augmented Generation (RAG) framework tailored for Arabic mental health Q&A. User queries are normalized and enhanced to handle dialectal variations, then matched with relevant contexts through hybrid retrieval, combining dense embeddings (Arabic-SBERT-100K) and sparse BM25 search. Retrieved contexts are re-ranked using semantic similarity, BM25 score, text length, and question similarity, with culturally sensitive filtering to ensure safe and appropriate advice. Finally, a fine-tuned Saka-14B model generates responses using prompts that integrate the user query, top contexts, domain-specific instructions, and cultural constraints. This approach allows RetAug to produce contextually relevant and culturally appropriate answers while effectively grounding the generation in retrieved knowledge.
@@ Acknowledgments
We thank the organizers, contributors, annotators, reviewers, and sponsors of AraHealthQA 2025. We extend our special gratitude to Walid Al-Eisawi for managing the Codabench platform for Track 2 and for drafting the corresponding sections of this manuscript.

### v3

@@ Abstract
We introduce AraHealthQA 2025, the Comprehensive Arabic Health Question Answering Shared Task, held in conjunction with ArabicNLP 2025 (co-located with EMNLP 2025). This shared task addresses the paucity of high-quality Arabic medical QA resources by offering two complementary tracks: MentalQA, focusing on Arabic mental health Q&A (e.g., anxiety, depression, stigma reduction), and MedArabiQ, covering broader medical domains such as internal medicine, pediatrics, and clinical decision making. Each track comprises multiple subtasks, evaluation datasets, and standardized metrics, facilitating fair benchmarking. The task was structured to promote modeling under realistic, multilingual, and culturally nuanced healthcare contexts. We outline the dataset creation, task design and evaluation framework, participation statistics, baseline systems, and summarize the overall outcomes. We conclude with reflections on the performance trends observed and prospects for future iterations in Arabic health QA [^1] Author order, excluding the first two lead authors, is alphabetical. The final author served in an advisory role..
AraHealthQA 2025:
The First Shared Task on Arabic Health Question Answering
Hassan Alhuzali,1 Walid Al-Eisawi ,2 Muhammad Abdul-Mageed,3 Chaimae Abouzahir2 Mouath Abu-Daoud,2 Ashwag Alasmari,4 Renad Al-Monef,4 Ali Alqahtani,4 Lama Ayash,4 Leen Kharouf,2 Farah E. Shamout,2 Nizar Habash2 1Umm Al-Qura University, 2New York University Abu Dhabi, 3The University of British Columbia, 4King Khalid University Correspondence: hrhuzali@uqu.edu.sa & farah.shamout@nyu.edu
@@ 2.1 Mental Health Benchmarks
To bridge this gap, the MentalQA dataset Alhuzali and Alasmari (2025); Alhuzali et al. (2024) was developed, providing annotated Arabic question-answer pairs that cover a variety of question types and answer strategies. This dataset supports the creation and evaluation of NLP systems capable of handling various mental health inquiries, forming the foundation of the Arabic Mental Health Question Answering Shared Task. Using the MentalQA dataset, this track provides a dedicated benchmark for Arabic mental health question-answering. It addresses the dual challenge of classification and response generation, creating a platform to systematically evaluate models in a culturally sensitive setting. Through this effort, MentalQA promotes research on the building of reliable and context-sensitive NLP systems for Arabic-speaking communities.
@@ 3.1.1 Subtask 1 and 2
For both subtasks, the dataset is based on MentalQA and is divided into 300 samples for training, 50 samples for development, and 150 samples as a blind test set for final evaluation. The training set can be used to fine-tune LLMs or serve as a base for few-shot learning approaches. The development set is intended to tune hyper-parameters and evaluate performance, while the test set ensures fair benchmarking of all participants.
@@ 4 Shared Task Teams
Submission Rules: For Track 1, we allowed participant teams to submit up to five runs for each test set and for each of the three subtasks. For Track 2, participants were initially allowed 10 submissions each, which was later increased to 15 submissions due to platform-specific issues. For each team, only the submission with the highest score was retained on the official leaderboard. The official evaluation relied on a blind test set. To ensure fairness and reproducibility, each subtask of each track was hosted as a separate competition on Codabench Xu et al. (2022), enabling automatic scoring and ranking of submissions. These Codabench instances will remain active even after the official competition concludes, supporting continued experimentation and benchmarking on the MentalQA and MedArabiQ datasets.
@@ 4.1 Track 1: MentalQA
Evaluation: Subtask 1 and Subtask 2 are multi-label classification tasks and are evaluated using Weighted F1 score and Jaccard score. The Weighted F1 balances precision and recall while accounting for class imbalance, whereas the Jaccard score measures the overlap between predicted and gold label sets, making it suitable for multi-label evaluation. Subtask 3 is evaluated using BERTScore Zhang et al. (2020), which leverages contextual embeddings from pre-trained language models to capture semantic similarity between generated responses and reference answers. Together, these metrics provide a robust assessment of system performance across the classification and generation subtasks, reflecting both the accuracy of the labels and the semantic quality of the outputs.
@@
MarsadLab bessghaier et al. (2025) Hamad Bin Khalifa University, Qatar; Northwestern University, Qatar 1, 2
Sindbad Morsy et al. (2025) George Washington University, USA 1, 2, 3
@@
RetAug AbdelAziz et al. (2025) Nile University, Egypt 1,2, 3
@@
Fahmni Sabty et al. (2025) MBZUAI, UAE; Gameball Company; German International 1, 2, 3
University, Egypt; American University in Cairo, Egypt
@@ 5.1.4 General Description of Submitted Systems (Track 1)
Subtask 1: The winning team, mucAI Abdou (2025), achieved a weighted F1-score of 0.61 for question classification. Their system, Explain–Retrieve–Verify (ERV), is a lightweight, training-free pipeline for multi-label categorization of Arabic mental-health questions. ERV combines a chain-of-thought LLM classifier with example-based retrieval and a verification agent. The LLM proposes candidate labels and rationales, a similarity agent retrieves top-$k$ nearest questions via multilingual sentence-transformer embeddings to provide case-based priors, and the verification agent reconciles these signals to produce a final label set with calibrated confidence. A post-processing step handles code parsing and confidence clamping. ERV runs efficiently at inference time without requiring fine-tuning or external data.
Subtask 2: The winning team, Sindbad Morsy et al. (2025), achieved a weighted F1-score of 0.71 and a Jaccard score of 0.71 for answer classification. Their approach leverages dataset augmentation to balance underrepresented classes, followed by a rigorous pipeline that uses state-of-the-art pre-trained language models (PLMs) and large language models (LLMs) for few-shot prompting and instruction fine-tuning. They utilize Gradient-free Edit-based Instruction Search (GrIPS) to optimize prompt selection, improving the quality and consistency of the QA system without extensive manual intervention.
Subtask 3: The winning team, RetAug AbdelAziz et al. (2025), achieved a BERTScore of 0.679 for generative question answering. Their system employs a Retrieval-Augmented Generation (RAG) framework tailored for Arabic mental health Q&A. User queries are normalized and enhanced to handle dialectal variations, then matched with relevant contexts through hybrid retrieval, combining dense embeddings (Arabic-SBERT-100K) and sparse BM25 search. Retrieved contexts are re-ranked using semantic similarity, BM25 score, text length, and question similarity, with culturally sensitive filtering to ensure safe and appropriate advice. Finally, a fine-tuned Saka-14B model generates responses using prompts that integrate the user query, top contexts, domain-specific instructions, and cultural constraints. This approach allows RetAug to produce contextually relevant and culturally appropriate answers while effectively grounding the generation in retrieved knowledge.
@@ Acknowledgments
We thank the organizers, contributors, annotators, reviewers, and sponsors of AraHealthQA 2025.